Solve each equation. identify each equation as a conditional equation, an inconsistent equation, or an identity.
1/2b - 1/2=1/4b

Answer by fractalier(1829)

Let us multiply all by 4b to clear the fractions, so we get
1/2b - 1/2 = 1/4b
4b[1/2b - 1/2 = 1/4b]
2 - 2b = 1
now subtract 2, then divide by -2 and get
-2b = -1
b = 1/2
Must be a conditional equation, since an inconsistent equation has no solution, and an identity has an infinite number of solutions...